OpenStudy (anonymous):

What is the mass of 12.0 ml of a substance with a density of 0.922 g/cm3? A. 11.1 B. 1.11 g C. 0.11 g D. 11.1 g

OpenStudy (anonymous):

d=m/v or v=m*d so v= 11.064, rounding it to 11.1 [A]<----